OpenStudy (anonymous):

Let C[−1,1] be the real inner product space consisting of all continuous functions f : [−1,1] → R, with the inner product := integral of f(X)g(x)dx from x=-1 to x=1 Let W be the subspace of odd functions, i.e. functions satisfying f(−x) = −f(x). Find (with proof) the orthogonal complement of W.

OpenStudy (anonymous):

@dan815 @oldrin.bataku

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

bout time

OpenStudy (anonymous):

i'm going to have dinner now. have a sus

OpenStudy (anonymous):

u need to multiply food by human to equal eat right now

OpenStudy (dan815):

is linear algebra fun?

OpenStudy (dan815):

does it expand the way you think about life?

OpenStudy (anonymous):

the orthogonal complement of W is the set of even functions

OpenStudy (anonymous):

here's the proof: we need to find\[g \in C[-1,1]\]such that \[<f,g> = 0\]\[\int\limits_{-1}^{1} f(x)g(x)\: dx=0\]\[\int\limits_{-1}^{0} f(x)g(x)\: dx + \int\limits_{0}^{1}f(x)g(x)\:dx = 0\]\[-\int\limits_{1}^{0}f(-x)g(-x) \: dx + \int\limits_{0}^{1} f(x)g(x)\: dx = 0\]\[\int\limits_{1}^{0}f(x)g(-x)\:dx+\int\limits_{0}^{1}f(x)g(x)\:dx=0\]\[-\int\limits_{0}^{1}f(x)g(-x)\:dx + \int\limits_{0}^{1}f(x)g(x)\:dx = 0\]\[\int\limits_{0}^{1}f(x)(g(x)-g(-x))\:dx=0\]it must be\[g(x) - g(-x) = 0\]\[g(x) = g(-x)\]it's clear that g is even

OpenStudy (anonymous):

i seee! interesting.

OpenStudy (anonymous):

oh ok. i see what u did with the integrals.

OpenStudy (anonymous):

be carfull, because this is the oposite thing. He proved (copy paste) complement for even functions, and you need to supose that f(x) is odd, :)

OpenStudy (anonymous):

the problem states that f(x) is odd

OpenStudy (anonymous):

the diff. is just the sign

OpenStudy (anonymous):

if f(x) is even, when you manipulate the integrals, you'll get g(x) + g(-x), and if f(x) is odd you'll get g(x) - g(-x)

OpenStudy (anonymous):

yea i had problems with manipulating the integrals...thanks for that.

OpenStudy (anonymous):

good like btw.

OpenStudy (anonymous):

good link*